### v3

# Absence of bulk charge density wave order in the normal state of UTe2

[…]
[name=W. Simeth, color=blue]WS
I Title
Absence of bulk charge density wave order in the normal state of UTe2
II Authors
C. S. Kengle1,2, J. Vonka3, S. Francoual4, J. Chang5, P. Abbamonte2, M. Janoschek5,6, P. F. S. Rosa1, and W. Simeth1,5,6
III Affiliations
1Los Alamos National Laboratory, Los Alamos, NM 87545, USA
2Department of Physics, University of Illinois Urbana-Champaign, Urbana, Illinois, 61801, USA
3Laboratory for X-ray Nanoscience and Technologies, Paul Scherrer Institute, Villigen PSI, Switzerland
4Deutsches Elektronen-Synchrotron (DESY), D-22607 Hamburg, Germany
5Physik-Institut, Universität Zürich, Winterthurerstrasse 190, CH-8057 Zürich, Switzerland
6Laboratory for Neutron and Muon Instrumentation, Paul Scherrer Institute, Villigen PSI, Switzerland
IV Abstract
A spatially modulated superconducting state, known as pair density wave (PDW), is a tantalizing state of matter with unique properties. Recent scanning tunneling microscopy (STM) studies revealed that spin-triplet superconductor UTe2 hosts an unprecedented spin-triplet, multi-component PDW whose three wavevectors are indistinguishable from a preceding charge-density wave (CDW) order that survives to temperatures well above the superconducting critical temperature, $T_{c}$. Whether the PDW is the mother or a subordinate order remains unsettled. Here, based on a systematic search for bulk charge order above $T_{c}$ using resonant elastic X-ray scattering (REXS), we show that the structure factor of charge order previously identified by STM is absent in the bulk within the sensitivity of REXS. Our results invite two scenarios: either the density-wave orders condense simultaneously at $T_{c}$ in the bulk, in which case PDW order is likely the mother phase, or the charge modulations are restricted to the surface.
V Introduction
Unconventional superconductors not only provide a platform for the investigation of exotic pairing mechanisms beyond electron-phonon coupling [1, 2, 3, 4], but already underpin existing technology such as powerful superconducting magnets or have potential for future applications such as topological quantum computing with fractionalized excitations [5]. In particular, most conventional $s$-wave superconductors are topologically trivial, whereas unconventional superconductors may be topological depending on the underlying superconducting (SC) order parameter. For any newly-discovered unconventional superconductor, the determination of its SC order parameter therefore becomes a central question. However, this task is rendered difficult due to experimental discrepancies, disorder, and intertwined orders, such as charge-density waves (CDWs), magnetism, pair-density waves (PDWs), and nematicity [6, 7].
UTe2 is a recent addition to the family of unconventional superconductors [8, 9], and its SC order parameter continues to defy consensus. Though there is strong evidence that UTe2 is a spin-triplet superconductor [10, 11, 12, 13], reports of chiral, multicomponent, and topological superconductivity [14, 15, 16, 17] have been challenged [18, 19, 20, 21, 22]. Recently, scanning tunneling microscopy (STM) measurements identified that the SC order parameter, $\Delta_{\bm{q}}(\bm{R})$, is spatially modulated at three wave-vector components, $\bm{q}_{i}$ ($i=$ 1, 2, 3), which are around the Brillouin zone boundary [23] and which display different field and temperature variations [23]. This pair-density wave (PDW) is inherently linked to CDW order, $\rho_{\bm{q}}(\bm{R})$, which displays the same modulation components as $\Delta_{\bm{q}}(\bm{R})$ but shifted by a phase difference $\pi$ [24, 23, 25].
Notably, CDW order survives to temperatures well above the superconducting critical temperature, $T_{c}\approx 1.6$ K. As discussed in Ref. [24], this intriguing result points to two possible scenarios of intertwined order: either the SC and PDW order parameters are dominant and generate subordinate charge modulations or the SC and CDW orders dominate and give rise to pair density modulations. In the first case, a superconducting gap energy of $250$ $\mu$eV coupled to a 10 $\mu$eV PDW state must induce a CDW with a much larger energy scale of order 25 meV. In the second case, the normal-state CDW coexists with superconductivity below $T_{c}$ to generate a 10 $\mu$eV PDW state at the same wavevectors. Gu et al [24] argue that the second case is more plausible, whereas Aishwarya et al [23] reason that a normal-state CDW is not inconsistent with a PDW mother phase because the PDW may melt into a CDW phase that survives to higher temperatures. Whether the PDW is the mother order or a subordinate order therefore remains unsettled.
Though STM is a powerful surface probe that provides both phase and domain sensitive information with high accuracy [26, 27], it is not inherently sensitive to ordered states in the bulk and therefore unable to make conclusive statements on the penetration depth of a state. In fact, signatures of CDWs observed in STM are equally likely to arise from bulk or surface states, and the two forms of order may coexist in a material [28]. Therefore, with all experimental evidence for either CDW or PDW modulations in UTe2 being restricted to surface sensitive measurements, it remains unclear whether these modulations in the SC state extend into the bulk and, if so, whether the broadening of the momentum space CDW peaks above $T_{c}$ is also a bulk phenomenon. Signatures of a phase transition when CDW order vanishes are absent in transport or thermodynamic measurements [29, 30, 31], therefore emphasizing the necessity to perform a systematic search for charge-density wave signatures using microscopic bulk-sensitive methods.
Here we provide the desired bulk measurements of the charge order in UTe2 via resonant elastic X-ray scattering (REXS) performed just above the superconducting transition ($T=2.2\,$ K), which is about four times lower than the highest temperature the CDW in UTe2 is reported to exist [25]. Resonant diffraction can increase intensity of charge order by more than three orders of magnitude compared to non-resonant scattering [32, 33], where intensities are frequently prohibitively weak and which is largely insensitive to weak forms of charge-order not involving lattice degrees of freedom. In particular, non-resonant scattering precludes to rule out the presence of bulk long-range order. Therefore, having used incident X-ray energies of 4.95 keV (Te $L_{\mathrm{1}}$ edge) and 3.73 keV (U $M_{4}$ edge) is central for the conclusion of our study on the absence of bulk charge order in UTe2.
Because STM above the superconducting transition identified nanometer-sized patches of charge order, we investigated a region in reciprocal space near two of the reported wave vectors with respect to broad signals of correlation lengths smaller than typical patch sizes. Additionally, to consider a state that has a different correlation length in the bulk, we investigated the vicinity of one of the ordering vectors on a tight grid sufficiently dense and large enough to identify resolution limited CDW-signals using incident X-ray energy of 3.73 keV. Our main result is that in both cases the structure factor of the putative CDW is absent from the bulk within our detection limits posed by resonant X-ray diffraction.
VI Results
VI.1 Surface charge order observed in STM
We first revisit the nature of charge order identified previously in STM studies [24, 23, 25]. Data were taken on $(011)$ planes of UTe2. The top layer after cleaving consists of Te atoms that form a 2-dimensional orthorhombic Bravais lattice with a rectangular centered unit cell (cf. Supplementary Materials Text 1 and Figure S1). Fig. 1(A) illustrates schematically characteristic signatures in two-dimensional reciprocal space, as seen in Fourier transformed STM images and discussed in Refs. [24, 23, 25]. The periodic arrangement of Te-atoms results in relatively strong structural Fourier-components (black circles). Below $T_{c}$ long-range charge-order is observed with components $\mathrm{q}_{\mathrm{1}}=(0,0.57)$, $\mathrm{q}_{\mathrm{2}}=(1,0.43)$, and $\mathrm{q}_{\mathrm{3}}=(-1,0.43)$, shown in Fig. 1(A) in terms of orange circles. With STM being restricted to atomic length scales and extending, at best, over two atomic layers, these studies were, however, unable to directly identify modulation components along $(011)$.
We will now consider putative bulk charge order with wave-vectors that result in these three surface projections. Although a single wave-vector (single-$\bm{q}$ with $\bm{q}=(0,0,0.57)$) lacking a modulation along $(011)$ could on its own lead to this pattern of projections, it is unable to account for the multi-component behavior observed through field and thermal variations (cf. Ref.[23]). We will therefore consider three independent components $\bm{q}_{\mathrm{1}}$, $\bm{q}_{\mathrm{2}}$, and $\bm{q}_{\mathrm{3}}$ in the three-dimensional Brillouin zone of UTe2.
[…]
Figure 1(B) illustrates a section of the 3D Brillouin zone for UTe2. The relevant characteristic high-symmetry points and estimated CDW wave vectors are show in purple and orange, respectively. Other high-symmetry points are indicated with grey spheres. The estimated CDW wave-vector positions do not match any obvious commensurate value, but are in the plane perpendicular to $\Gamma$S ($\Gamma$ is the origin). The first wave-vector, $\bm{q}_{\mathrm{1}}=(0.57,0.5,0.5)$, is slightly off the W point, $\bm{q}_{\mathrm{2}}=(0.430,0.338,1.339)$ and $\bm{q}_{\mathrm{3}}=(0.430,0.662,-0.339)$ are close to the Brillouin zone corners L1 and L2.
[…]
Above the superconducting transition temperature the charge order begins to melt, but persists in the form of patches which shrink upon increasing temperature [34]. For example, at 4.8 K the patches are of the order of order $5$ nm in size. Finally, we use exemplary STM data above $T_{c}$ to infer a conservative estimation of the expected full width at half maximum (FWHM) of bulk CDW order. Fig. 1(C) shows a typical intensity $vs.$ momentum curve obtained from a Fast Fourier Transformation (FFT) of STM images taken at 4.9 K along the line connecting the points W and S, $i.e.$, between $(\frac{1}{2},\frac{1}{2},\frac{1}{2})$ and $(\frac{1}{2},\frac{1}{2},0)$ as specified in conventional orthorhombic coordinates. A pronounced peak near the W point arises from charge-density wave component $\bm{q}_{1}$ and is characterized by a broad Gaussian profile centered at $H=0.545\,$ r.l.u..
Consider now, that the peak FWHM seen in STM represents a convolution of intrinsic peak shape ($f_{0}=2/\kappa_{0}$), associated with the correlation length $\kappa_{0}$ of CDW order, and the experimental resolution of the STM instrument, ($f_{\mathrm{STM}}$). Assuming that both $f_{0}$ and $f_{\mathrm{STM}}$ describe the widths of Gaussian profiles, we find that the profile in Fig. 1(B) has a FWHM $f_{\mathrm{W,STM}}=\sqrt{f_{0}^{2}+f_{\mathrm{STM}}^{2}}$. The experimental profile therefore serves as an upper bound for the internal width of charge-order, yielding $f_{\mathrm{max}}:=f_{\mathrm{W,STM}}=0.1439\,\mathrm{\AA}^{-1}>f_{0}$. In turn, a lower bound may be obtained considering that coherent CDW correlations were observed in terms of patches of less than $\kappa_{1}=5$ nm diameter, implying the bound for the intrinsic peak width of $f_{\mathrm{min}}:=2/\kappa_{1}<f_{0}$, cf. Ref. [35]. Combined, these limits on the FWHM can be written $f_{\mathrm{min}}=0.04\,\mathrm{\AA}^{-1}<f_{0}<0.1439\,\mathrm{\AA}^{-1}=f_{\mathrm{max}}$.
Figure 2: Resolution-limited structural Bragg peaks. (A) At incident energy of $3.73$ keV, the FWHM is $4.6985\times 10^{-4}$ r.l.u. = $6.85\times 10^{-4}\,\mathrm{\AA}^{-1}$ (B) At incident energy of $4.95$ keV, the FWHM is $3.07\times 10^{-4}$ r.l.u. $=4.68\times 10^{-4}\mathrm{\AA}^{-1}$.
VI.2 Resonant X-ray scattering
To study bulk microscopic properties of charge modulations in UTe2, we utilized resonant elastic scattering of hard, linearly polarized X-rays and searched for signatures of a multi-component CDW around the Brillouin zone boundary. With penetration depths in UTe2 exceeding 200200200\,200nm, these experiments are sensitive to the structure factor of long-range correlations across several hundred crystallographic unit cells.
For the putative resonant enhancement of charge-density wave order, the precise structure factor is unknown. But it is well established, that in the non-resonant case when using linearly polarized X-rays diffraction intensities from charge-order may have maximum intensity in the polarization channel $\sigma\sigma^{\prime}$, i.e., with both incident and scattered polarization parallel to the scattering plane, [36]. Therefore, in order to maximize the chance to observe charged-ordered Bragg peaks we maintain incident polarization parallel to the scattering plane [36].
In our experiments, structural Bragg peaks measured on UTe2 were limited by experimental resolution represented by Gaussian profiles, indicating ideal crystalline quality of our sample. Figure 2 shows representative scans through (A) the $(011)$ Bragg peak using an incident energy of $3.73$ keV and (B) the $(022)$ Bragg peak using an incident energy of $4.95$ keV. See Supplementary Information, Text S2, for more details on the setup of our diffraction experiments.
VI.3 Survey for putative CDW order in bulk
To search for the CDW in the bulk, we will consider two limits. First, that the CDW peaks are resolution limited (having the same resolution as the structural Bragg peaks). Next, that the CDW has correlation lengths of order $5$ nm, inspired by previous STM measurements [25].
Figure 3: Absence of resolution-limited CDW Bragg peaks near $\bm{q}_{\mathrm{1}}$. (A) Reciprocal space region that was investigated by means of $H$-scans. $K_{0}$ and $L_{0}$ coordinates were held constant during each scan and are illustrated in terms of colored spheres. The grey sphere denotes the W-point, the projection where the CDW-peak is expected according to STM. The green ellipsoid illustrates the resolution of the diffractometer. (B) $H$ scan at $(K,L)=(2.4996,2.4856)$, zoomed in to emphasize the shape of a resolution-limited Gaussian of height $3\sigma$. (C) REXS intensity of all $H$-scans. Colors denote the $K_{0}$, $L_{0}$-coordinates in (A). Plots are offset for clarity, and each horizontal line indicates zero for the corresponding plot. The intensity scale of 2.5 cts/s is indicated by the black scale bar.
We start with the search for resolution-limited charge-order peaks. Such long-range order has correlation lengths much larger than $2/f_{\mathrm{res}}$, where $f_{\mathrm{res}}$ denotes the width in momentum space of resolution-limited structural Bragg peaks. To increase our chance to discover charge-order and to avoid missing its signatures due to the high resolution of REXS, we chose an incident energy of 3.37 keV, where the resolution volume is larger compared to 4.95 keV.
We surveyed reciprocal space in the search of signal at $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.54,2.5,2.5)$, corresponding to CDW wave-vector of type $\bm{q}_{1}$, shown by the grey sphere in the center of Fig. 3(A). The letter $a$ denotes the specific momentum transfer, where the CDW peak was studied in our experiments. To account for possible uncertainties in momentum transfers due to the accuracy of the diffractometer (see Supplementary Information Text S3), we performed a 3-dimensional grid search at positions near $\bm{Q}^{\mathrm{CDW}}_{1,a}$. The mesh consists of $H$-scans at given $K_{0}$, $L_{0}$-coordinates indicated in Fig. 3 in terms of spheres. The spacing along $K$ and $L$ was chosen as 0.0018 r.l.u apart, which is smaller than the instrument resolution (green ellipsoid in Fig. 3(A)). The step width in each linescan along $H$, $\Delta H=3.6\cdot 10^{-4}\,\mathrm{\AA}^{-1}$, is smaller by a factor of two than the experimental resolution determined for a nearby specular Bragg peak (c.f. Supplementary Material Text 3), and therefore small enough to account for uncertainties in the incommensurate component of $\bm{q}_{1}$ along $(100)$.
[…]
Figure 4: Absence of diffuse charge-scattering peaks in resonant X-ray diffraction data. The three panels show momentum-space cuts through positions, where according to STM reports diffuse charge scattering peaks of a multi-component CDW above the superconducting transition were reported. Shown in (A) is an $H$ scan through $\bm{q}_{\mathrm{1}}$ at ($H,2.5,2.5$) using 3.73 keV incident energy, in (B) through $\bm{q}_{\mathrm{1}}$ at ($H,3.5,3.5$) using 4.95 keV incident energy, and in ((C)) through $\bm{q}_{\mathrm{2}}$ at ($H,2.338,2.661$) using incident energy 4.95 keV. As explained in the text and derived from STM, a diffuse CDW signal may be expected to have full width have maximums in the interval $f_{\mathrm{min}}=0.04\,\mathrm{\AA}^{-1}<f_{0}<0.1439\,\mathrm{\AA}^{-1}=f_{\mathrm{max}}$. The red circles present X-ray counts, where the error bars correspond to uncertainties arising from Poisson counting statistics. Points at zero intensity are here shown with vanishing error bars. The background was fitted by linear functions (blue thick line). The green line corresponds to a Gaussian peak with an amplitude of $3\sigma$ above the background, where $\sigma$ denotes the mean error bar of the X-ray counts.
Turning now to the second scenario, where CDW peaks have $\sim$ nm-sized short correlation lengths, we expand our search to look for signal at the corresponding bulk positions of $\bm{q}_{\mathrm{1}}$ and $\bm{q}_{\mathrm{2}}$. For wavevector $\bm{q}_{\mathrm{1}}$, data were collected in two Brillouin zones using two incident energies: $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.54,2.5,2.5)$ at 3.73 keV and $\bm{Q}^{\mathrm{CDW}}_{1,b}=(-0.54,3.5,3.5)$ at 4.95 keV. Here $b$ labels another momentum-transfer, where we studied the respective wave-vector. For wavevector $\bm{q}_{\mathrm{2}}$, scans were taken in one Brillouin zone at $\bm{Q}^{\mathrm{CDW}}_{2,b}=(-0.430,2.338,2.661)$ with incident energy of 4.95 keV.
The results of these scans for both incident energies and wavevectors are presented in Fig. 4. Panels (A) and (B) show scans at $\bm{q}_{\mathrm{1}}$ at 3.73 keV and 4.95 keV, respectively, and panel (C) shows scans at $\bm{q}_{\mathrm{2}}$ at 4.95 keV. The background intensities were each fitted using a line with a constant slope and offset, shown in blue. The fit results are displayed above each panel. Mean error bars, $\sigma$, as obtained from Poisson counting statistics are given by (A) $0.17\,\mathrm{cts}/\mathrm{s}$, (B) $8.9\,\mathrm{cts}/\mathrm{s}$, and (C) $8.38\,\mathrm{cts}/\mathrm{s}$.
A charge ordered peak having a FWHM of $(f_{\mathrm{min}}+f_{\mathrm{max}})/2$ and amplitude $\geq 3\sigma$ would be detectable above the background, as indicated by the green Gaussian curves in Fig. 4. We conclude that such Gaussian profiles are absent within our detection limit in all three cases.
Comparison of intensities lets us conclude that the maximum intensities of CDW peaks having short correlation lengths in bulk on the Te L-edge are at least eight orders of magnitude smaller than for the structural Bragg peaks. Compared to that, in the cuprate YBa2Cu3O6.67 [37] even for non-resonant x-rays charge order signals are only seven orders of magnitude smaller relative to strong nearby structural Bragg peaks.
VI.4 Upper boundary on atomic displacements
Charge modulations probed by Thomson scattering of X-rays can arise either from spatial modulation of valence electrons or from periodic atomic displacements [38]. These two effects typically appear together and cause each other. With periodic lattice displacements leading to a much stronger response in X-ray scattering, we search for them directly and are further interested in them as indirect evidence of valence modulations [37]. Here, we will calculate the limits of atomic displacements in UTe2 based on the scattering intensities reported above.
Using standard expressions, harmonic displacements of atoms along a direction $\bm{d}$ away from the equilibrium position $\bm{R}_{i}$ may be written as $\bm{R}^{\prime}_{i}=\bm{R}_{i}+u\bm{\hat{d}}\cdot\sin(\bm{q}_{0}\cdot\bm{R}_{i})$. In diffraction, this leads to charge-order Bragg peaks around structural Bragg peaks $\bm{G}$, appearing in leading order at $\bm{G}\pm\bm{q}_{0}$.
Neglecting higher order terms, the structure factor of the modulated crystal around $\bm{Q}=\bm{G}$ can be written as
$S_{\mathrm{m}}(\bm{Q})=S(\bm{G})\delta(\bm{Q}-\bm{G})+S(\bm{G})\biggl{(}\frac{\bm{Q}\cdot\bm{u}}{2}\biggr{)}^{2}\delta(\bm{Q}-\bm{G}\pm\bm{q}_{0})\,.$ (1)
As the structure factor is proportional to the integrated intensities of Bragg peaks, i.e., $I(\bm{Q})\propto S(\bm{Q})$, the ratio of intensities of a structural Bragg peak to a charge order Bragg peak therefore yields for the amplitude $\bm{u}=u\bm{d}$ of the modulation around $\bm{G}$
$\frac{I(\bm{q}_{0}+\bm{G})}{I(\bm{G})}=\biggl{(}\frac{\bm{u}\cdot(\bm{q}_{0}+\bm{G})}{2}\biggr{)}^{2}$ (2)
The intensity at a given position in momentum space may arise from a combination of different charge-order satellites associated with neighboring Brillouin zones. For example, at momentum $\bm{Q}^{\mathrm{CDW}}_{1,a}=(-0.545,2.5,2.5)$, the intensity arises from CDW modulations with wave-vector component $(-0.545,0.5,0.5)$ of the $\bm{G}_{1}=(0,2,2)$ zone and with $(-0.545,-0.5,-0.5)$ of the $\bm{G}_{2}=(0,3,3)$ zone, two different representatives of the $\bm{q}_{1}$-modulation. Calculation of the crystal structure factor of UTe2 (see Supplementary Material Text S5) permits comparison with the intensity of the structural Bragg peak at $\bm{G}_{0}=(0,1,1)$ and yields
$\left(\frac{\bm{u}\cdot\bm{Q}^{\mathrm{CDW}}_{1,a}}{2}\right)^{2}=\frac{A(\bm{Q}^{\mathrm{CDW}}_{1,a})\cdot V(\bm{Q}^{\mathrm{CDW}}_{1,a})}{A(\bm{G}_{0})\cdot V(\bm{G}_{0})}\cdot\frac{S(\bm{G}_{0})}{S(\bm{G}_{1})+S(\bm{G}_{2})}.$ (3)
Following this procedure, we first look at the measurements taken at 3.37 keV shown in Sec. VI.3. For resolution-limited CDW peaks maximum atomic displacements can be estimated using the above equation for $V(\bm{Q}^{\mathrm{CDW}}_{1,a})\approx V(\bm{G}_{0})$. Along the (100) direction we determine $u_{||(100)}\leq 4.7\cdot 10^{-3}\,\mathrm{\AA}$ and along the (110) direction $u_{||(011)}\leq 1.2\cdot 10^{-3}\,\mathrm{\AA}$. In comparison, the atomic displacements in the CDW in YBa2Cu3O6.67 are of order $u\approx 3\cdot 10^{-3}\,\mathrm{\AA}$ [37], or similarly $u\approx 10^{-3}\,\mathrm{\AA}$ in related compounds [39].
Now we turn to the data shown in Fig. 4, where we searched for long-range CDW order with nanometer-sized correlation lengths. Again using Eq. 3 with consideration of multiple domains, we can find upper bounds for displacement amplitudes of CDWs having widths $f_{\mathrm{CDW}}=(f_{\mathrm{min}}+f_{\mathrm{max}})/2=0.09\,\mathrm{\AA}^{-1}$. Since the CDW peak width is much larger than the REXS momentum resolution, the integration volume is essentially resolution independent: $V(\bm{Q}_{\mathrm{CDW}})=(f_{\mathrm{CDW}}^{2}+f_{\mathrm{res}}^{2})^{3/2}\approx f_{\mathrm{CDW}}^{3}$. Considering data from Fig. 4(B), we find that atomic displacements at 4.95 keV incident energy are limited by $u\leq 0.59\,\mathrm{\AA}$ along the $(100)$ direction and $u\leq 0.1\,\mathrm{\AA}$ along the $(011)$ direction.
While our data set relatively strong lower bounds on absolute intensity limits of charge order with short correlation lengths, the bounds on maximum displacement amplitudes are still relatively weak. The reason here is that short correlation lengths naturally extend over larger regions in momentum space and result, in turn, in relatively small diffraction peak amplitudes ($V(\bm{Q}^{\mathrm{CDW}}_{1,a})\ll V(\bm{G}_{0})$ in Eq. 3). Therefore, we conclude that finite displacements with low correlation lengths are not ruled out by our data.
VII Discussion
Our measurements using resonant X-rays show that the charge modulations earlier observed in STM above the superconducting transition at the surface of UTe2 are absent from the bulk within our detection limits posed by resonant X-ray diffraction. Our measurements were performed at $T=2.2$ K, about four times lower than the highest temperature the CDW in UTe2 is reported to exist [25]. Our results did not reveal diffraction intensity in the vicinity of the CDW wave vectors at either the U M4 or Te L1 absorption edges above the background level.
The coverage of our fine mesh-grid search near the $\bm{q}_{1}$ CDW wave vector rules out the presence of sharp resolution-limited charge-order peaks. Notably, the maximum diffraction intensity of such a CDW peak at the U M edge is at least five orders of magnitude smaller relative to the intensity of the strongest Bragg peaks observed. Atomic displacements of such modulations are four orders of magnitude smaller than the respective atomic units along $(100)$ and $(011)$.
We further investigated reciprocal space with respect to CDW peaks having short correlation lengths in the bulk. Comparing REXS intensities, we find that maximum intensities of such charge order peaks are at least eight orders of magnitude weaker than nearby structural Bragg peaks. We found further that finite atomic displacements of such modulations are not ruled out by our data.
We highlight that neither thermodynamic measurements [29, 30, 31] nor inelastic neutron scattering identified any signatures suggestive of bulk charge order above the superconducting transition [40, 41, 42, 43]. In combination with our results, the emergence of a surface charge-density wave in the normal state provides the most consistent picture. As for the superconducting state, our results point to two possible conclusions: either the density-wave orders condense simultaneously at $T_{c}$ in the bulk, in which case PDW order is likely the mother phase, or these charge modulations are restricted to the surface. Bulk scattering measurements in the superconducting state are required to address this outstanding question.
We note that two additional experimental investigations of UTe2 at low temperature–a non-resonant X-ray diffraction measurement [44] as well as pulse echo and resonant ultrasound measurements [45]–have been reported concurrently with this work. Notably, both studies, performed independently and on samples grown by different groups, are in agreement with our results. Our study is complementary and using resonant X-rays we provide bounds on the structure factor posed detection limits in our experiments.
VIII Materials and Methods
Sample Preparation
[…]
Resonant Elastic X-ray Scattering
Experiments were carried out in the second experimental hutch EH2 of beamline P09 at the synchrotron source PETRA III [46]. The X-ray diffraction was carried out in a horizontal scattering geometry. The UTe2 sample was cooled using a variable temperature insert in a cryomagnet able to provide magnetic fields up to 14 T. X-rays with incident linear polarization were chosen using phase plates [47].
We use the conventional orthorhombic basis with lattice parameters $a=4.16\,\mathrm{\AA}$, $b=6.13\,\mathrm{\AA}$, and $c=13.97\,\mathrm{\AA}$ for the description of UTe2 crystals. Momentum transfers are denoted by capital bold italic letters, $\bm{Q}$, and given in the reference-frame of the three-dimensional Brillouin zone of UTe2 by means of $\bm{Q}=H\bm{b}_{1}+K\bm{b}_{2}+L\bf{b}_{3}$, where $\bm{b}_{i}=\epsilon_{ijk}2\pi/V\cdot\bm{b}_{j}\times\bm{b}_{k}$. $H$, $K$, $L$ are the Miller indices and $\epsilon_{ijk}$ is the Levi-Civita symbol and $V$ the volume of the conventional unit cell in real-space. Propagation vectors of bulk long-range order are denoted by lowercase bold italic letters, $\bm{q}$. Momentum transfers of structural Bragg peaks are labelled with the letter $\bm{G}$
Characteristic high-symmetry points in the three-dimensional Brillouin zone of UTe2 are denoted by capital letters. Points and coordinates in the two-dimensional reciprocal lattice, used to describe the Fourier transformation of STM images, are given in lowercase letters.
Structural Bragg peaks measured on UTe2 essentially featured Gaussian profiles. Their intensity may be modelled by:
$G(\bm{Q}):=A\cdot E(\bm{Q})\,$ (4)
where $A$ denotes the amplitude (or the maximum) and $E$ is a Gaussian profile with full width at half maxima $f_{\bm{d}_{1}}$, $f_{\bm{d}_{2}}$, and $f_{\bm{d}_{3}}$ along the three directions $\bm{d}_{1}=(1,0,0)$, $\bm{d}_{2}=(0,1,1)$, and $\bm{d}_{3}=\bm{d}_{1}\times\bm{d}_{2}$, respectively. Integrated intensity of a Bragg peak is therefore given by $I=\frac{1}{8}\sqrt{\frac{\pi}{\ln(2)}}^{3}\cdot Af_{\bm{d}_{1}}f_{\bm{d}_{2}}f_{\bm{d}_{2}}$. The profile of structural Bragg peaks is essentially restricted by resolution. We therefore define $f_{\bm{d}_{1}}f_{\bm{d}_{2}}f_{\bm{d}_{2}}$ as the experimental resolution volume in momentum space.
Structure factors of UTe2 for X-ray scattering were calculated with scattering amplitudes provided by the International Tables for Crystallography, Ref [32]. Details on these calculations are provided in the Supplementary Information, Text S5.
Data Availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on request.
Acknowledgments
WS acknowledges fruitful discussions with Aline Ramirez. The authors further thank Simon Gerber for discussions. WS was supported through funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Sklodowska-Curie grant agreement No 884104 (PSI-FELLOW-III-3i). Work at Los Alamos National Laboratory was performed under the auspices of the U.S. Department of Energy, Office of Basic Energy Sciences, Division of Materials Science and Engineering. C.S.K. acknowledges support from the Laboratory Directed Research and Development program. P.A. acknowledges support from the Gordon and Betty Moore Foundation, EPiQS grant GBMF9452. J. C. acknowledges support from Swiss National Science foundation under grant 200021-188564. MJ acknowledges funding by the Swiss National Science Foundation through the project ”Berry-Phase Tuning in Heavy f-Electron Metals” (200650). We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III at DESY. Beamtime was allocated for proposal I-20221340 EC. Author contributions: MJ, PFSR, and WS conceived the study. MJ and WS designed the experiments. PFSR grew the samples. CK, JV, SF, and WS carried out the X-ray experiments. CK, PFSR and WS interpreted the results. CK, PFSR, and WS wrote the paper with input from all the authors. Competing interests: The authors declare that they have no competing interests. Data and materials availability: All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. Correspondence and requests for additional materials may be addressed to wsimeth@lanl.gov.